OpenStudy (youngnephew34):

7x+7=4y Can anybody help me PLZZZZZZZZZZZZZZZZZZZ I WILL GIVE YOU A MEDAL AND FAN

OpenStudy (anonymous):

Is the y supposed to be by the 4?

OpenStudy (youngnephew34):

yes

OpenStudy (youngnephew34):

Can anybody else help me

OpenStudy (misty1212):

HI!!

OpenStudy (youngnephew34):

@EmilyF03 @plzzhelpme @Tazmaniadevil

OpenStudy (misty1212):

you wrote \[7x+7=4y\] but you did not ask a question

OpenStudy (youngnephew34):

It says what is the vaule of x in this diagram

OpenStudy (anonymous):

Um which variable are u trying to isolate ?

OpenStudy (misty1212):

\(x\) is a variable it can be anything

OpenStudy (misty1212):

if you want to solve \[7x+7=4y\] for \(x\) your answer will have a \(y\) in it \[7x+7=4y\\ 7x=4y-7\\ x=\frac{4y-7}{7}\]

OpenStudy (youngnephew34):

but I'm trying to figure it out and it gave me 10

OpenStudy (misty1212):

can you post a screenshot or something? i am thinking there is some typo here

OpenStudy (youngnephew34):